OpenStudy (noahbred):

Please help me? Lanie presses the gas pedal in her car down so she can pass another car. She was going 60 mph but ends up going 80 mph in order to pass another car. It takes her car 5 seconds to make this increase in speed. What is Lanie’s acceleration during this time? 12 mph/second 28 mph/second 16 mph/second 4 mph/second

OpenStudy (noahbred):

so 25

OpenStudy (noahbred):

please tell me its correct

OpenStudy (whimsical):

\[80-60= x\] \[5-0=y\] \[\frac{ x }{ y }= acceleration\] find x and y first

OpenStudy (noahbred):

ok si divide both

OpenStudy (anonymous):